OpenStudy (anonymous):

The cost to produce football uniforms for a school can be modeled by C(u) = 1/4 u2 - 25u +3500 where C(u) represents the cost in dollars to produce u football uniforms. a. Find the cost to produce 30 uniforms. b. Find the vertex and describe the meaning c. How many uniforms can a school get with a budget of $1600?

OpenStudy (anonymous):

1) replace \(u\) by \(30\) , that is, compute \[C(30)=\frac{1}{4}\times 30^2-25\times 30+3500\]

OpenStudy (anonymous):

2) first coordinate of the vertex of \(ax^2+bx+c\) is always \(-\frac{b}{2a}\) in your example \(a=\frac{1}{4},b=-25\). compute \(-\frac{-25}{2\times \frac{1}{4}}=25\times 2=50\) to find the second coordinate, find \(C(50)\)

OpenStudy (anonymous):

this tells you that the minimum occurs when they produce 50 uniforms (seems rather unlikely, they really stretch for these word problems, don't they?)

OpenStudy (anonymous):

3) set \[\frac{1}{4}u^2-25u+3500=1600\] and solve for \(u\) but there is no real solution , so either there is a typo in the question, or it is a trick question and the answer is "they can't buy any"